# Rua Dr. Roberto Frias, 4200-465 Porto, Portugal

### v1

@@ 1 Introduction
Motivated by this observation, the first goal of this work is to provide machine-translated versions of FairytaleQA, in three of the most widely spoken Romance languages: Spanish, Portuguese (European pt-PT and Brazilian pt-BR), and French. To estimate the quality of the translated versions, we conduct an error analysis of one of the translated datasets. As our second goal, we establish baseline benchmarks for both QA and QG tasks, achieved through fine-tuning modest-scale models. These models offer the advantage of being usable on more affordable hardware, in contrast to large language models that might incur costs for using expensive hardware or external APIs. Figure 1 shows an example where the original text and a QA pair are translated into Portuguese and applied in both QA and QG tasks.
@@ 5.2 Evaluation Protocol
To evaluate the quality of the generated QA pairs, we engaged a group of 15 volunteers who are native speakers of Portuguese with a higher education degree. Each individual was instructed to read a real-exam narrative text and rate a set of 15 mixed QA pairs about that text. The set included 5 QA pairs directly extracted from a real exam, 5 generated by the QAPG model [^10] The 5 questions were randomly selected from a pool of 7, corresponding to the 7 narrative elements that can be controlled., and 5 generated by the GPT-4111111We prompt GPT-4 Turbo to generate QA pairs from the narrative text in the target language, tailored for children aged 7 to 10, aligned with the FairytaleQA’s audience. model. We conducted 4 sets of inquiries, totaling 60 QA pairs evaluated. The texts were sourced from real exams provided by the Portuguese Ministry of Education [^12] https://iave.pt/provas-e-exames/provas-e-exames/provas-de-afericao-eb/ for elementary school students, aligning with FairytaleQA’s target audience. For each question, participants were instructed to assess it based on the following metrics:

### v2

@@ 1 Introduction
Motivated by this observation, the first goal of this work is to provide machine-translated versions of FairytaleQA, in three [^2] We also have included translated datasets for Italian and Romanian in our repository, although they were not studied in this research. of the most widely spoken Romance languages: Spanish, Portuguese (European pt-PT and Brazilian pt-BR), and French. To estimate the quality of the translated versions, we conduct an error analysis of one of the translated datasets. As our second goal, we establish baseline benchmarks for both QA and QG tasks, achieved through fine-tuning modest-scale models. These models offer the advantage of being usable on more affordable hardware, in contrast to large language models that might incur costs for using expensive hardware or external APIs. Figure 1 shows an example where the original text and a QA pair are translated into Portuguese and applied in both QA and QG tasks.
@@ 5.2 Evaluation Protocol
To evaluate the quality of the generated QA pairs, we engaged a group of 15 volunteers who are native speakers of Portuguese with a higher education degree. Each individual was instructed to read a real-exam narrative text and rate a set of 15 mixed QA pairs about that text. The set included 5 QA pairs directly extracted from a real exam, 5 generated by the QAPG model [^11] The 5 questions were randomly selected from a pool of 7, corresponding to the 7 narrative elements that can be controlled., and 5 generated by the GPT-4121212We prompt GPT-4 Turbo to generate QA pairs from the narrative text in the target language, tailored for children aged 7 to 10, aligned with the FairytaleQA’s audience. model. We conducted 4 sets of inquiries, totaling 60 QA pairs evaluated. The texts were sourced from real exams provided by the Portuguese Ministry of Education [^13] https://iave.pt/provas-e-exames/provas-e-exames/provas-de-afericao-eb/ for elementary school students, aligning with FairytaleQA’s target audience. For each question, participants were instructed to assess it based on the following metrics: